OpenStudy (anonymous):

I need the slope for (4,3) an (-5,-2) please. Someone help I will help work it out

OpenStudy (anonymous):

I know that 4 and -5 are x and 3 and -2 are y

OpenStudy (anonymous):

m = change y/change x m = 3+2/4+5 = 5/9

OpenStudy (anonymous):

how is this possible

OpenStudy (anonymous):

I see that you have added them but it just does not seem right

myininaya (myininaya):

its the formula for slope slope is rise/run=change of y/change of x=(y1-y2)/(x1-x2)

myininaya (myininaya):

m=(-2-3)/(-5-4)=-5/-9=5/9

myininaya (myininaya):

or the same thing the other way m=(3-(-2))/(4-(-5))=(3+2)/(4+5)=5/9

OpenStudy (anonymous):

ok 100% sure please and thank you I am trying to learn these so that PI doesnt think I am on here just for the answer.